# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 1. Introduction
with $\operatorname{\mathcal{O}}_{F_{1}}/(p)\simeq\kappa_{1}[T]/(T^{e})=\kappa_{1}[\varepsilon_{1}]$ action, such that $\kappa_{1}$ acts on $\mathscr{F}_{j}^{l}$ by $\sigma_{j}:\mathcal{O}_{F_{1}^{\rm ur}}\to\kappa$; for all $1\leq l\leq e$, $\mathscr{F}_{j}^{l}/\mathscr{F}_{j}^{l-1}$ is locally free of rank $d_{j}^{l}$ and finally $\varepsilon_{1}\mathscr{F}_{j}^{l}\subset\mathscr{F}_{j}^{l-1}$. The $\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}$-action on $\Lambda$ induces a decomposition $\Lambda=\bigoplus_{j}\Lambda_{j}$. For each embedding $\sigma^{l}_{j}:\operatorname{\mathcal{O}}_{F_{1}}\hookrightarrow\operatorname{\mathcal{O}}_{F}$ which extends $\sigma_{j}:\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}\hookrightarrow\operatorname{\mathcal{O}}_{F}$, we set $\Lambda_{j}^{l}=\Lambda_{j}\otimes_{\operatorname{\mathcal{O}}_{F_{1}},\sigma^{l}_{j}}\operatorname{\mathcal{O}}_{F}$ (note that the lattice $\Lambda_{j}^{l}$ is denoted by $\Xi_{j}^{l}$ in [47]*Proposition 5.2) and $\Lambda^{l}=\bigoplus_{j}\Lambda_{j}^{l}$. Set $\Lambda_{j,0}^{l}=\Lambda_{j}^{l}\otimes_{\operatorname{\mathcal{O}}_{F}}\kappa$ and $\Lambda^{\mathrm{spl}}_{0}=\bigoplus_{l}\bigoplus_{j}\Lambda_{j,0}^{l}$. One can show that there is a standard $F$-zip structure on $\Lambda^{\mathrm{spl}}_{0}$, whose Hodge filtration is given by the cocharacter $\mu$ over $\kappa$.
For $(\underline{{A}},\underline{\mathscr{F}_{\bullet}})\in{\mathscr{A}_{0}}^{\mathrm{spl}}(S)$, one can associate a module
@@
Then each $\mathcal{M}_{j}^{l}$ is a locally free $\operatorname{\mathcal{O}}_{S}$-module, and locally isomorphic to $\Lambda_{j}^{l}\otimes_{\operatorname{\mathcal{O}}_{F}}\operatorname{\mathcal{O}}_{S}$, cf. [47]. Now the idea is to transfer the Hodge and conjugate filtrations on $\mathcal{H}$ to filtrations on each $\mathcal{M}^{l}:=\bigoplus_{j}\mathcal{M}_{j}^{l}$, and to show that the $F$-zip structure on $\mathcal{H}$ induces an $F$-zip structure on $\mathcal{M}$. To this end, one can first complete the filtration of $\omega$ to a full filtration of $\mathcal{H}$, then apply the Frobenius and Verschiebung morphisms successively to the full filtration, and make use of the relation $\textrm{Ker}\,F=\textrm{Im}\,V,\textrm{Ker}\,V=\textrm{Im}\,F$ to succeed. In fact, we proceed by an alternative and equivalent approach. For any $1\leq j\leq f$, inspired by [54] we construct explicit linear maps $F_{j}^{l}:\mathcal{M}_{j}^{l-1}\to\mathcal{M}_{j}^{l},\;V_{j}^{l}:\mathcal{M}_{j}^{l}\to\mathcal{M}_{j}^{l-1}$ for each $2\leq l\leq e$ (which are easy to construct), and $\sigma^{-1}$-linear ($\sigma$-linear) maps $F_{j}^{1}:\mathcal{M}_{j-1}^{e}\to\mathcal{M}_{j}^{1},\;V_{j}^{1}:\mathcal{M}_{j}^{1}\to\mathcal{M}_{j-1}^{e}$ (which are induced by the Frobenius and Verschiebung, and the definition of $V_{j}^{1}$ is a little subtle). Now for each $1\leq l\leq e$, $\mathcal{M}^{l}=\bigoplus_{j=1}^{f}\mathcal{M}_{j}^{l}$ is a locally free $\operatorname{\mathcal{O}}_{S}[\varepsilon_{1}]$-module with $\kappa_{1}$ action, and locally $\operatorname{\mathcal{M}}^{l}\simeq\Lambda^{l}\otimes_{\operatorname{\mathcal{O}}_{F}}\operatorname{\mathcal{O}}_{S}.$ For each $l$, set $F^{l}=\bigoplus_{j}F_{j}^{l}$ and $V^{l}=\bigoplus_{j}V_{j}^{l}$. We have the following linear morphisms:
@@
[…]
@@
    There is a canonical isomorphism $g:\operatorname{\mathcal{M}}^{1}\to\operatorname{\mathcal{M}}^{1}$ sends $\operatorname{Ker}\,V^{1}$ to $\operatorname{Im}\,F^{1}$.
@@
over $\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}$. Moreover, there are natural compatible symplectic form and $\operatorname{\mathcal{O}}_{L}$-action on $\mathcal{M}$. Comparing with the standard $F$-zip structure on $\Lambda^{\mathrm{spl}}_{0}$, we get the $\mathcal{G}^{\mathrm{spl}}_{0}$-zip of type $\mu$ over $\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}$. Roughly, the process of transforming $\mathcal{H}$ to $\operatorname{\mathcal{M}}$ is a way of “semi-simplification”. During this transformation the different types of $\mathcal{G}^{\operatorname{rdt}}_{0}$-zips become uniform as $\mathcal{G}^{\mathrm{spl}}_{0}$-zips. For more details, see Proposition 3.8 and subsection 3.2.
@@
which we call the Ekedahl-Oort (EO) stratification. Here the index set ${}^{J}W$ is the subset of the (absolute) Weyl group of $\mathcal{G}^{\mathrm{spl}}_{0}$ defined by the cocharacter $\mu$ in the usual way, which is equipped with the partially order $\preceq$ as that in [51, 52]. By the theorem, each stratum $\operatorname{\mathscr{A}}^{\mathrm{spl},w}_{k}$ is non-empty, smooth, and we have the usual closure relation for the EO strata given by $\preceq$. In fact, the non-emptiness is more subtle, requiring additional efforts. We prove the non-emptiness by comparing the minimal EO stratum of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ with the minimal EKOR stratum of $\operatorname{\mathscr{A}}_{k}$ (see below) in Proposition 3.11, and show the minimal EKOR stratum is non-empty in the Appendix (Proposition A.15) by adapting the method of He-Zhou [24].
@@
  • Kottwitz-Rapoport (KR) stratification $\operatorname{\mathscr{A}}_{k}=\coprod_{w\in\textrm{Adm}(\mu)_{K}}\operatorname{\mathscr{A}}^{w}_{k}$, which is induced from the local model diagram for $\operatorname{\mathscr{A}}_{k}$ and the geometry of $\mathbb{M}^{\mathrm{loc}}_{k}$, cf. [47, 38, 19].
  • Ekedahl-Kottwitz-Oort-Rapoport (EKOR) stratification $\operatorname{\mathscr{A}}_{k}=\coprod_{x\in{}^{K}\textrm{Adm}(\mu)}\operatorname{\mathscr{A}}^{x}_{k}$, which can be constructed as in [57], see the Appendix of the current paper; in particular, see Example A.11 for more information of EKOR strata in the Hilbert case.
@@
The Newton stratification on $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ naturally factors through $\operatorname{\mathscr{A}}_{k}$. In subsection 3.6 we will show that the Hodge stratification also factors through $\operatorname{\mathscr{A}}_{k}$. Moreover, the resulted stratification on $\operatorname{\mathscr{A}}_{k}$ is coarser than the Kottwitz-Rapoport stratification. On the other hand, the Ekedahl-Oort stratification on $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ does not factor through $\operatorname{\mathscr{A}}_{k}$ in general. Nevertheless, one can try to compare the EO strata of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ with pullbacks of KR and EKOR strata of $\operatorname{\mathscr{A}}_{k}$ under $\pi$. On each KR stratum $\operatorname{\mathscr{A}}_{k}^{w}$, we have a $\mathcal{G}_{0}^{\textrm{rdt}}$-zip of type $J_{w}$, see [57] and our Appendix. Pulling back to $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$, we get a family of $\mathcal{G}_{0}^{\textrm{rdt}}$-zips of different types. It is curious to study the relationship with our $\mathcal{G}^{\mathrm{spl}}_{0}$-zip of fixed type $\mu$. In doing so, we prove the following (see Propositions 3.15 and 3.16 in subsection 3.5)
  • Let $\operatorname{\mathscr{A}}_{k}^{w_{0}}$ be the maximal KR stratum of $\operatorname{\mathscr{A}}_{k}$. Then $\pi$ induces an isomorphism $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\operatorname{\mathscr{A}}_{k}^{w_{0}}$; furthermore $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})$ is also the maximal Hodge stratum (called the generalized Rapoport locus in [4]).
  • $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})$ is a disjoint union of some EO strata of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$, and moreover
  • these EO strata are exactly pullbacks of the EKOR strata contained in $\operatorname{\mathscr{A}}_{k}^{w_{0}}$, thus the isomorphism $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\operatorname{\mathscr{A}}_{k}^{w_{0}}$ refines into isomorphisms between the corresponding EO and EKOR strata.
In particular, we deduce that the maximal EO stratum coincides with the $\mu$-ordinary locus. Thus by Theorem 1.1 we reprove the open density of $\mu$-ordinary locus of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$, cf. [4].
@@
we get Hasse invariants on $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$ by pulling back the group theoretic Hasse invariants on the stack $\mathcal{G}^{\mathrm{spl}}_{0}\textrm{-Zip}^{\mu}_{\kappa}$ constructed by Goldring-Koskivirta in [18]. Note that in [3] Bijakowski and Hernandez have constructed Hasse invariants for the $\mu$-ordinary locus (which is also our maximal EO stratum) by an explicit method. It would be interesting to compare their construction with our construction here.
@@
Theorems 1.1 and 1.2 allow us to deduce some further consequences on the mod $p$ geometry and arithmetic related to PEL-type Shimura varieties with ramifications raised from Weil restrictions. As an illustration, we discuss congruences of mod $p$ automophic forms and Galois representations associated to torsion classes in coherent cohomology. For some other possible applications, see [12] and [63]. From Theorem 1.2, we get extensions of Hasse invariants to toroidal compactifications. Then we can apply Goldring-Koskivirta’s machinery of Hasse-regular sequences introduced in [18], which formalizes some key properties of the subschemes defined by Hasse invariants. Indeed, the key technical conditions 6.4.2 and 7.1.2 of [18] have been verified in our setting as above. Let $S$ be the finite set of primes $\ell$ where $K_{\ell}$ is not hyperspecial. Let $L\subset P_{\mu}$ be the Levi subgroup and $(\eta,V)\in\mathrm{Rep}_{F}L$ an irreducible representation of highest weight $\eta$. Then we get the associated integral automorphic vector bundle $\mathcal{V}_{\eta}$ on $\mathsf{M}_{K}^{\mathrm{spl}}$ and the subcanonical extension $\mathcal{V}_{\eta}^{\mathrm{sub}}$ of $\mathcal{V}_{\eta}$ to $\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$. For each integers $i\geq 0,n\geq 1$, the Hecke algebra $\mathcal{H}^{S}$ acts on the cohomology groups
@@
where $\varpi$ is a uniformizer of $\operatorname{\mathcal{O}}_{F}$. Let $\mathcal{H}^{i,n}_{\eta}$ be its image in the endomorphism algebra. Fix a representation $r:{}^{L}G\rightarrow\mathrm{GL}_{m}$ of the Langlands dual group. With all the geometric ingredients at hand, one can prove the following theorem by the same arguments of [18].
Theorem 1.3.
@@
For unramified PEL type case, Theorem 1.3 is due to [18] (the key technical conditions 6.4.2 and 7.1.2 there hold for unramified PEL type case), which recovers the main result of [14] (for the unramified Hilbert case), and completes the work of Boxer [5] by a different approach. We refer to [18] subsection 10.1 for further discussions on the appearance of non-degenerate limit of discrete series in higher coherent cohomology of Shimura varieties. Recall that since $K_{p}\subset G(\mathbb{Q}_{p})$ is a very special parahoric subgroup, irreducible smooth representations $\pi_{p}$ of $G(\mathbb{Q}_{p})$ such that $\pi^{K_{p}}_{p}\neq 0$ can be classified by their spherical parameters, see [68] section 6. The interests of Theorem 1.3 lie in that it provides some flexibilities for the construction of automorphic Galois representations, in the sense we allow the $p$-component of the automorphic representation $\pi$ to be ramified as $\pi^{K_{p}}_{p}\neq 0$. It reproves [54] Theorem 1.1 in the Hilbert case by a different method. We also mention that a similar result has been given in [50] in the ramified setting, but there the integral structure is given by the strange integral models introduced by Scholze using perfectoid geometry. As all these works, Theorem 1.3 should be useful when studying modularity lifting problems, cf. [8].
@@
Acknowledgments. We thank Sian Nie for some helpful discussions. The first author is partially supported by the National Key R&\&&D Program of China 2020YFA0712600, the CAS Project for Young Scientists in Basic Research, Grant No. YSBR-033, and the NSFC grant No. 12288201.

### v3

@@ 1. Introduction
with $\operatorname{\mathcal{O}}_{F_{1}}/(p)\simeq\kappa_{1}[T]/(T^{e})=\kappa_{1}[\varepsilon_{1}]$ action, such that $\kappa_{1}$ acts on $\mathscr{F}_{j}^{l}$ by $\sigma_{j}:\mathcal{O}_{F_{1}^{\rm ur}}\to\kappa$; for all $1\leq l\leq e$, $\mathscr{F}_{j}^{l}/\mathscr{F}_{j}^{l-1}$ is locally free of rank $d_{j}^{l}$ and $\varepsilon_{1}\mathscr{F}_{j}^{l}\subset\mathscr{F}_{j}^{l-1}$.
The $\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}$-action on $\Lambda$ induces a decomposition $\Lambda=\bigoplus_{j}\Lambda_{j}$. For each embedding $\sigma^{l}_{j}:\operatorname{\mathcal{O}}_{F_{1}}\hookrightarrow\operatorname{\mathcal{O}}_{F}$ which extends $\sigma_{j}:\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}\hookrightarrow\operatorname{\mathcal{O}}_{F}$, we set $\Lambda_{j}^{l}=\Lambda_{j}\otimes_{\operatorname{\mathcal{O}}_{F_{1}},\sigma^{l}_{j}}\operatorname{\mathcal{O}}_{F}$ (note that the lattice $\Lambda_{j}^{l}$ is denoted by $\Xi_{j}^{l}$ in [49]*Proposition 5.2) and $\Lambda^{l}=\bigoplus_{j}\Lambda_{j}^{l}$. Set $\Lambda_{j,0}^{l}=\Lambda_{j}^{l}\otimes_{\operatorname{\mathcal{O}}_{F}}\kappa$ and $\Lambda^{\mathrm{spl}}_{0}=\bigoplus_{l}\bigoplus_{j}\Lambda_{j,0}^{l}$. One can show that there is a standard $F$-zip structure on $\Lambda^{\mathrm{spl}}_{0}$, whose Hodge filtration is given by the cocharacter $\mu$ over $\kappa$.
For $S={\mathscr{A}_{0}}^{\mathrm{spl}}$ and $(\underline{{A}},\underline{\mathscr{F}_{\bullet}})\in{\mathscr{A}_{0}}^{\mathrm{spl}}(S)$ the universal object, one can associate a module
@@
Then each $\mathcal{M}_{j}^{l}$ is a locally free $\operatorname{\mathcal{O}}_{S}$-module, and locally isomorphic to $\Lambda_{j}^{l}\otimes_{\operatorname{\mathcal{O}}_{F}}\operatorname{\mathcal{O}}_{S}$, cf. [49]. Now the idea is to transfer the Hodge and conjugate filtrations on $\mathcal{H}$ to filtrations on each $\mathcal{M}^{l}:=\bigoplus_{j}\mathcal{M}_{j}^{l}$, and to show that the $F$-zip structure on $\mathcal{H}$ induces an $F$-zip structure on $\mathcal{M}$. To this end, one can first complete the filtration of $\omega$ to a full filtration of $\mathcal{H}$, then apply the Frobenius and Verschiebung morphisms successively to the full filtration, and make use of the usual relation $\textrm{Ker}\,F=\textrm{Im}\,V,\textrm{Ker}\,V=\textrm{Im}\,F$ to succeed. In fact, we proceed by an alternative and equivalent approach. For any $1\leq j\leq f$, inspired by [56] we construct explicit linear maps $F_{j}^{l}:\mathcal{M}_{j}^{l-1}\to\mathcal{M}_{j}^{l},\;V_{j}^{l}:\mathcal{M}_{j}^{l}\to\mathcal{M}_{j}^{l-1}$ for each $2\leq l\leq e$ (which are easy to construct), and $\sigma^{-1}$-linear ($\sigma$-linear) maps $F_{j}^{1}:\mathcal{M}_{j-1}^{e}\to\mathcal{M}_{j}^{1},\;V_{j}^{1}:\mathcal{M}_{j}^{1}\to\mathcal{M}_{j-1}^{e}$ (which are induced by the Frobenius and Verschiebung, and the definition of $V_{j}^{1}$ is a little subtle). Now for each $1\leq l\leq e$, $\mathcal{M}^{l}=\bigoplus_{j=1}^{f}\mathcal{M}_{j}^{l}$ is a locally free $\operatorname{\mathcal{O}}_{S}[\varepsilon_{1}]$-module with $\kappa_{1}$ action, and locally $\operatorname{\mathcal{M}}^{l}\simeq\Lambda^{l}\otimes_{\operatorname{\mathcal{O}}_{F}}\operatorname{\mathcal{O}}_{S}.$ For each $l$, set $F^{l}=\bigoplus_{j}F_{j}^{l}$ and $V^{l}=\bigoplus_{j}V_{j}^{l}$. We have the following linear morphisms:
@@
${{\mathcal{M}^{e}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F^{1}}$${{\mathcal{M}^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{V^{1}}$
@@
    There is a canonical isomorphism $g:\operatorname{\mathcal{M}}^{1}\to\operatorname{\mathcal{M}}^{1}$ sending $\operatorname{Ker}\,V^{1}$ to $\operatorname{Im}\,F^{1}$.
@@
[…]
@@
which we call the Ekedahl-Oort (EO) stratification. Here the index set $^{J}W$ is the subset of the (absolute) Weyl group of $\mathcal{G}^{\mathrm{spl}}_{0}$ defined by the cocharacter $\mu$ in the usual way, which is equipped with the partially order $\preceq$ as that in [53, 54]. By Theorem 1.1, each stratum $\operatorname{\mathscr{A}}^{\mathrm{spl},w}_{k}$ is non-empty, smooth, and we have the usual closure relation for the EO strata given by $\preceq$. In fact, the non-emptiness is more subtle, requiring additional efforts. We prove the non-emptiness by comparing the minimal EO stratum of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ with the minimal EKOR stratum of $\operatorname{\mathscr{A}}_{k}$ (see below) in Proposition 3.11, and show the minimal EKOR stratum is non-empty in the Appendix (Proposition A.15) by adapting the method of He-Zhou [26].
@@
  • Kottwitz-Rapoport (KR) stratification $\operatorname{\mathscr{A}}_{k}=\coprod_{w\in\textrm{Adm}(\mu)_{K_{p}}}\operatorname{\mathscr{A}}^{w}_{k}$, which is induced from the local model diagram for $\operatorname{\mathscr{A}}_{k}$ and the geometry of $\mathbb{M}^{\mathrm{loc}}_{k}$, cf. [49, 40, 21].
  • Ekedahl-Kottwitz-Oort-Rapoport (EKOR) stratification $\operatorname{\mathscr{A}}_{k}=\coprod_{x\in{}^{K_{p}}\textrm{Adm}(\mu)}\operatorname{\mathscr{A}}^{x}_{k}$, which can be constructed as in [59], see the Appendix of the current paper; in particular, see Example A.11 for more information on EKOR strata in the Hilbert case.
@@
The Newton stratification on $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ naturally factors through $\operatorname{\mathscr{A}}_{k}$. In subsection 3.6 we will show that the Hodge stratification also factors through $\operatorname{\mathscr{A}}_{k}$. Moreover, the resulted stratification on $\operatorname{\mathscr{A}}_{k}$ is coarser than the Kottwitz-Rapoport stratification. On the other hand, the Ekedahl-Oort stratification on $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ does not factor through $\operatorname{\mathscr{A}}_{k}$ in general. Nevertheless, one can try to compare the EO strata of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$ with pullbacks of KR and EKOR strata of $\operatorname{\mathscr{A}}_{k}$ under $\pi$. On each KR stratum $\operatorname{\mathscr{A}}_{k}^{w}$, we have a $\mathcal{G}_{0}^{\textrm{rdt}}$-zip of type $J_{w}$, see [59] and our Appendix. Pulling back to $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$, we get a family of $\mathcal{G}_{0}^{\textrm{rdt}}$-zips of different types. It is curious to study the relationship with our $\mathcal{G}^{\mathrm{spl}}_{0}$-zip of fixed type $\mu$. In doing so, we prove the following results.
Theorem 1.2 (Propositions 3.15 and 3.16).
  1. (1)
    Let $\operatorname{\mathscr{A}}_{k}^{w_{0}}$ be the maximal KR stratum of $\operatorname{\mathscr{A}}_{k}$. Then $\pi$ induces an isomorphism
    $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\operatorname{\mathscr{A}}_{k}^{w_{0}};$
    furthermore $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})$ is also the maximal Hodge stratum (called the generalized Rapoport locus in [6]).
  2. (2)
    $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})$ is a disjoint union of some EO strata of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$, and moreover these EO strata are exactly pullbacks of the EKOR strata contained in $\operatorname{\mathscr{A}}_{k}^{w_{0}}$. Thus the isomorphism $\pi^{-1}(\operatorname{\mathscr{A}}_{k}^{w_{0}})\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\operatorname{\mathscr{A}}_{k}^{w_{0}}$ refines into isomorphisms between the corresponding EO and EKOR strata.
In particular, using a result of He-Nie [24], we can deduce that the maximal EO stratum coincides with the $\mu$-ordinary locus, cf. Corollary 3.17. Thus by Theorem 1.1 we reprove the open density of $\mu$-ordinary locus of $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{k}$, cf. [6].
@@
we get Hasse invariants on $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$ by pulling back the group theoretic Hasse invariants on the stack $\mathcal{G}^{\mathrm{spl}}_{0}\textrm{-Zip}^{\mu}_{\kappa}$ constructed by Goldring-Koskivirta in [20]. Note that in [5] Bijakowski and Hernandez have constructed Hasse invariants for the $\mu$-ordinary locus (which is also our maximal EO stratum) by an explicit method. It would be interesting to compare their construction with our construction here. On the other hand, in the Hilbert case, in subsection 4.4 we do compare the EO strata and Hasse invariants here with those constructed in [56].
@@
Theorems 1.1 and 1.3 allow us to deduce some further consequences on the mod $p$ geometry and arithmetic related to PEL-type Shimura varieties with ramifications arising from Weil restrictions. As an illustration, we discuss congruences of mod $p$ automophic forms and Galois representations associated to torsion classes in coherent cohomology. For some other possible applications, see [14] and [66]. From Theorem 1.3, we get extensions of Hasse invariants to toroidal compactifications. Then we can apply Goldring-Koskivirta’s machinery of Hasse-regular sequences introduced in [20], which formalizes some key properties of the subschemes defined by Hasse invariants. Indeed, the key technical conditions 6.4.2 and 7.1.2 of [20] have been verified in our setting as above. Let $S$ be the finite set of primes $\ell$ where $K_{\ell}$ is not hyperspecial. Let $P_{\mu}\subset G_{F}$ be the parabolic subgroup associated to the cocharacter $\mu:\mathbb{G}_{m,F}\rightarrow G_{F}$ in the usual way (see the paragraph above Corollary 2.4), $\mathbf{L}\subset P_{\mu}$ the Levi subgroup and $V_{\eta}\in\mathrm{Rep}_{F}\mathbf{L}$ an irreducible representation of highest weight $\eta$. Using our modified local model diagram, we get the associated integral automorphic vector bundle $\mathcal{V}_{\eta}$ on $\mathsf{M}_{K}^{\mathrm{spl}}$ and the subcanonical extension $\mathcal{V}_{\eta}^{\mathrm{sub}}$ of $\mathcal{V}_{\eta}$ to $\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$. For each integers $i\geq 0,n\geq 1$, the Hecke algebra $\mathcal{H}^{S}$ acts on the cohomology groups
@@
where $\varpi$ is a uniformizer of $\operatorname{\mathcal{O}}_{F}$. Let $\mathcal{H}^{i,n}_{\eta}$ be its image in the endomorphism algebra. Fix a representation $r:{}^{L}G\rightarrow\mathrm{GL}_{m}$ of the Langlands dual group over $\mathbb{C}$ (and fix an isomorphism $\mathbb{C}\simeq\overline{\mathbb{Q}}_{p}$). With all the geometric ingredients at hand, one can prove the following theorem by the same arguments as [20].
Theorem 1.4 (Theorem 6.3).
@@
For the unramified PEL type case, Theorem 1.4 is due to [20] (the key technical conditions 6.4.2 and 7.1.2 there hold for the unramified PEL type case), which recovers the main result of [16] (for the unramified Hilbert case), and completes the work of Boxer [7] by a different approach. We refer to [20] subsection 10.1 for further discussions on the appearance of non-degenerate limit of discrete series in higher coherent cohomology of Shimura varieties. Recall that since $K_{p}\subset G(\mathbb{Q}_{p})$ is a very special parahoric subgroup, irreducible smooth representations $\pi_{p}$ of $G(\mathbb{Q}_{p})$ such that $\pi^{K_{p}}_{p}\neq 0$ can be classified by their spherical parameters, see [71] section 6. The significance of Theorem 1.4 lie in the flexibility it provides for the construction of automorphic Galois representations, in the sense that we allow the $p$-component of the automorphic representation $\pi$ to be ramified as $\pi^{K_{p}}_{p}\neq 0$. It reproves [56] Theorem 1.1 in the Hilbert case by a different method. We also mention that a similar result has been given in [52] in the ramified setting, but there the integral structure is given by the strange integral models introduced by Scholze using perfectoid geometry. As all these works demonstrate, Theorem 1.4 should be useful when studying modularity lifting problems, cf. [10].
In this paper we restrict to the PEL case, since so far splitting models for Shimura varieties have only been constructed in this setting ([49]). It would be interesting and useful to construct splitting models in more general Hodge and abelian type cases. This would certainly require quite different ideas and independent treatments. Once more general splitting models are available, we do expect that our constructions in this paper would naturally extend.
@@
Acknowledgments. We thank Sian Nie for some helpful discussions. We would like to thank the referee for careful reading and detailed comments, which help to improve the paper.